OpenStudy (anonymous):

Find three solutions of the equation y = 9x – 4. (1 point) • (–5, –49), (–2, –22), (3, 23) • (–5, –49), (2, –22), (3, 23) • (–5, –49), (–2, 22), (–3, 23) • (5, –49), (–2, 22), (–3, 23)

OpenStudy (luffingsails):

Hi @farmgirl11, the way I see to proceed is trial and error. Substitute in an x value from your points and see if the result is the correspond y value.

OpenStudy (anonymous):

what?

OpenStudy (luffingsails):

so, you are given: y = 9x-4 your first point is (-5, -49) So, in your equation, let x = -5 y = 9(-5) - 4 = -45 - 4 = -49 So, your first point (-5,-49) is a correct point for the function. See?

OpenStudy (anonymous):

not really most all of them says (-5,-49)

OpenStudy (luffingsails):

Right, but you need to test the other points in the set to see if they work. Some of them will fail.

OpenStudy (anonymous):

Its B

OpenStudy (anonymous):

@luffingsails is it B

OpenStudy (luffingsails):

Sorry, I keep getting dropped. Hold one sec.

OpenStudy (luffingsails):

B is right!

OpenStudy (anonymous):

thanks

OpenStudy (luffingsails):

You're welcome.